OpenStudy (anonymous):

the area of a rectangle is 9x^2-9x-4 cm^2. find its length and width then solve the perimeter Am I missing something here? The roots are -1/3 and 4/3 but I can't seem to figure out how they would come out as the length & width PLEASE HELP

OpenStudy (amoodarya):

\[ 9x^2-9x-4=(3x+1)(3x-4)\\area=\frac{1}{2} base * height =(3x+1)(3x-4)\\\]can you go on ?

OpenStudy (anonymous):

Yeah, I'm a bit stuck in this part so the length = (3x+1) width = (3x-4) ?

OpenStudy (anonymous):

or should I just leave it that way, like I don't have to find the exact value? :)

OpenStudy (amoodarya):

you missed 1/2

OpenStudy (amoodarya):

base *height =2 (3x+1)(3x-4) =(6x+2)(3x-4) or (3x+1) (6x-8)

OpenStudy (amoodarya):

base =(3x+1) height=6x-8 and viceversa base =(6x+2) height=3x-4 and viceversa

OpenStudy (anonymous):

but I thought the are of a rectangle is just l x w? :)

OpenStudy (anonymous):

*area

OpenStudy (amoodarya):

oh ...i am sorry i make it for a triangle but it is easier now let's do it again h*w=(3x+1)(3x-4) now we have to find 2(h+w) 2((3x+1)+(3x-4))

OpenStudy (anonymous):

Ohh, okay. So those are the answers? I'll just leave it with all the x's right? :)

OpenStudy (anonymous):

or is there a way we could find the exact values? :((

OpenStudy (amoodarya):

is there any information more ? if not , your question is complete

OpenStudy (anonymous):

no, nothing else. thank you so much! :)